# Dynamical Constraints on the Vertical Structure of Jupiter’s Polar Cyclones

### v1

@@ 3 Estimating the Deformation Radius Using a Single-Layer QG Model of the Jovian Poles
Figure 3: Interpretation of the Cyclones’ Vertical Structure from the Deformation Radius. The first row corresponds to the "mode 0" solutions of the eigenvalue problem described by Eq. 6, while the second row corresponds to "mode 1". (a,d) Vertical profiles of $\Phi_{n}$ for each mode. (b,e) Solutions for $L_{d}$ (green shades) for various combinations of $N$ (Brunt-Väisälä frequency) and $H$ (depth). The estimated $L_{d}$ distributions from Fig. 2k are shown by the blue and orange curves for the north and south poles, respectively. (c,f) Solutions for $\Phi_{n}$ as a function of pressure below the cloud layer, following the $N$ and $H$ relationship indicated by the blue solid curve in panels (b) and (e). Green dotted lines indicate the maximum sensitivity depths of the MWR channels. The numbers on the right ordinate show the altitude relative to 1 bar. Grey dashed lines in panels (b,c,e,f) denote the values used to generate the profiles in panels (a) and (d), and to produce Fig. 4. The color along the grey dashed line in panels (c,f) corresponds to the $\Phi(z)$ profiles shown in panels (a,d).
@@
4 Interpreting the Vertical Structure of Jupiter’s Polar Atmosphere from the Drift Rate of the Polar Cyclones
@@
The standard shallow-water (SW) expression for the deformation radius, $L_{d}=\sqrt{gH}/f_{0}$, is not well-suited to Jupiter’s upper atmosphere. This expression assumes a constant density and no vertical dependence of the flow, whereas Jupiter’s upper-atmospheric density increases exponentially with depth Guillot and Morel (1995) over many scale heights. To account for these vertical variations, we consider the full QG vorticity equation Vallis (2017):
@@
where $N$ is the Brunt-Väisälä frequency, indicating the stability of the atmosphere with respect to vertical motion, and $\tilde{\rho}$ is the density profile.
@@
where $\Phi_{n}(z)$ is the vertical profile (or "mode") for each $n$. Substituting this form into the QG equations requires each $\Phi_{n}$ to satisfy the vertical eigenvalue problem:
@@
where $\Gamma_{n}$ is an eigenvalue. For each mode $n$, the corresponding deformation radius is given by $L_{d,(n)}=\Gamma_{n}^{-\frac{1}{2}}$. Under the assumption that the cyclones are dominated by a single vertical mode, the horizontal dynamics derived from the two-dimensional QG equation (Eq. 3), and the full three-dimensional QG equation (Eq. 4) will yield the same horizontal dynamics for the given $L_{d,(n)}$. Given our estimation of the $L_{d}$ value required to achieve the observed westward drift (Fig. 3), we can use the eigenvalue problem posed by Eq. 6 to investigate which vertical structures (eigenfunctions $\Phi_{n}(z)$) satisfy the required $L_{d}$.
@@
We solve this eigenvalue problem using the Dedalus library Burns et al. (2020), applying boundary conditions of no vertical velocity at the tropopause $(\partial_{z}\Phi_{n}|_{{}_{z=0}}=0)$ and no flow at the bottom $(\Phi_{n}|_{{}_{z=-H}}=0)$. For $\tilde{\rho}(z)$, we adopt a reference density profile estimated by internal models of Jupiter, constrained by an AI-based parameter sweep Ziv et al. (2024a, b). We focus here on the lowest two modes ($n=\{0,1\}$), which are illustrated in Fig. 3(a,d). The $n=0$ mode corresponds to a simple decay profile, whereas $n=1$ changes signs along the column.
@@
Since Eq. 6 depends only on $N$ (assumed constant here) and $H$, we present $L_{d,(n)}$ solutions as a function of these two parameters in Fig. 3(b,e), for both modes. The constrained range of deformation radii from the previous section (Fig. 2) is indicated for the north (blue) and south (orange) poles, essentially giving an $H$ as a function of $N$ for each pole. Focusing on the relationship defined by the blue solid curve in Fig. 3(b,e), we plot the corresponding $\Phi_{n}(z)$ profiles as a function of pressure below the cloud level for different $N$ values (Fig. 3(c,f)).
@@
The MWR instrument, expected to probe below Jupiter’s clouds at the north pole in the near future, measures brightness temperature at six channels. The maximum sensitivity pressure levels for each MWR channel are marked by green dotted lines, with Channel 6 above the 1-bar level. For a plausible range $N=3-20\times 10^{-3}\rm{s}^{-1}$ Lee and Kaspi (2021), and focusing first on mode 0, we anticipate that the polar cyclones’ footprint will be visible in Channels 6 and 5 (and Channel 4 if $N\approx 0.3\times 10^{-2}\rm{s}^{-1}$). However, if the real Jovian poles exhibit turbulence with a vertical phase shift—reminiscent of terrestrial storms with low-level convergence and upper-level divergence—then mode 1 may also be excited. In that scenario, signatures could extend into Channel 3 for the same range of $N$s.
@@
Figure 4 illustrates an example of how $\psi$ extends downward and intersects with the MWR channels for $N=3\times 10^{-3}\,\mathrm{s}^{-1}$, the lower bound reported in Lee and Kaspi (2021). Channel 6, sensitive to pressures around $0.7$ bar, lies slightly above the eigenfunctions’ range (which starts near 1 bar). We assume that Channel 6 will display a $\psi$ field similar to that of Channel 5. The JIRAM background photograph Adriani et al. (2018) represents the cloud deck at approximately this pressure level. The downward extension of $\psi$ below Channel 4 depends on the excitation of mode 1. It is important to note that the MWR measures brightness temperature, which is only indirectly related to $\psi$. Predicting MWR observations requires a layer of translation, involving solutions for vertical velocity and the subsequent diffusion-advection of ammonia, given the background ammonia distribution Duer et al. (2021).
@@
Figure 4: Vertical Structure and Intersections with Instrument Sensitivity Depths at the North Pole. The $\psi$ field from Fig. 2c was extended downward using the eigenfunctions from Fig. 3a,d, which correspond to $N=3\times 10^{-3}\;(\mathrm{s}^{-1})$ and $H=41.7$ ($112.7$) km for mode 0 (mode 1). Both modes yield $L_{d}=192$ km, with mode 00 terminating ($\psi=0$) at $z=-41.7$ km, and mode 1 continuing further to $\psi=0$ at $z=-112.7$ km. The planes represent $\psi$ at the MWR channels that overlap with the eigenfunction range, while the shaded regions between the planes depict $\psi$ values between the channels. The dashed black line indicates the north pole. The background image on channel 6 is a JIRAM measurement from PJ4 Adriani et al. (2018), similar to Fig. 2a.
5 Discussion
@@
We then employed a 2D polar model initialized with identical cyclones at their observed locations (during PJ4). This model provides a dynamical constraint on the deformation radius, $L_{d}$, which encodes information about the cyclones’ vertical structure. We found that both excessively large and excessively small values of $L_{d}$ produce westward drifts that deviate strongly from observations. A sampling algorithm estimates $L_{d}\sim 200\,(\mathrm{km})$ at the north pole and $\sim 300\,(\mathrm{km})$ at the south pole. While the posterior distributions of $L_{d}$ at the two poles do overlap, it remains an open question whether genuine differences in $L_{d}$ exist between the poles, given Jupiter’s negligible obliquity and the resulting expected near-symmetry. However, the indicated difference in $L_{d}$ between the poles could have important implications for the cyclones’ characteristics. Specifically, the smaller $L_{d}$ at the north pole might place a constraint on the maximum possible size of the north polar cyclones, leading to their smaller size and, consequently, a larger number of cyclones.
Using these estimates of $L_{d}$, we sought consistency between the 2D QG framework (Fig. 2) and a 3D QG model that incorporates density variations with depth (Eq. 4). This required solving the eigenvalue problem in Eq. 6, in which the eigenvalues yield $L_{d}$ for different vertical modes, and the numerically determined eigenfunctions describe the cyclone’s vertical structure relative to the cloud deck. In our idealized approach, we assumed a constant Brunt-Väisälä frequency, $N$, and a single dominant vertical mode, both of which can introduce some uncertainty. The solutions (Fig. 3) suggest that for $N\approx 10^{-2}\,\mathrm{s}^{-1}$, cyclone depths of $10-80\,\mathrm{km}$ at the north pole and $20-120\,\mathrm{km}$ at the south pole are plausible, depending on the mode. If $N\approx 10^{-3}\,\mathrm{s}^{-1}$, the estimated cyclone depths increase substantially, ranging $100-450\,\mathrm{km}$ at the north pole and $200-600\,\mathrm{km}$ at the south pole.
@@
In summary, our work offers a novel way to bridge observed cloud-level motions and the deeper structure of Jovian polar cyclones. The constraints we derive on the deformation radius, vertical structure, and static stability at Jupiter’s poles can guide future studies aimed at modeling the energy fluxes responsible for the formation and maintenance of these cyclones against dissipation. By linking the deformation radius derived from $\beta$-drift constraints to vertical modes in a 3D QG framework, we gain a clearer understanding of how these cyclones might extend below the visible cloud layer, and how their vertical structure governs their horizontal drift.
@@
The data analyzed in this study were published by Ref.Mura et al. (2022) (DOI: https://doi.org/10.1029/2022JE007241), as cited in the text. The simulations were run with the Dedalus solver Burns et al. (2020) (DOI: https://doi.org/10.1103/PhysRevResearch.2.023068).
@@ Acknowledgments
This research has been supported by the Helen Kimmel Center for Planetary Science at the Weizmann Institute of Science. We thank Rei Chemke, Keren Duer, Eli Galanti and Or Hadas for insightful conversations and helpful feedback.
@@
Supporting Information
Developing a Model Following a Cyclone That Includes the Effect of Stretching
@@
$\frac{D}{Dt}\left(\xi+\beta y-\frac{1}{L_{d}^{2}}\psi\right)=0,$ (S1)
@@
$\frac{D}{Dt}=\left(\frac{\partial}{\partial t}+\mathbf{u}\cdot\nabla\right),$ (S2)
$\mathbf{u}$ is the velocity vector, $\xi=\nabla\times\mathbf{u}$ is the relative vorticity, and $\beta$ is the constant meridional gradient of the planetary vorticity ($f$) or any prescribed background vorticity gradient.
@@
$\begin{array}[]{c}\left(x,y\right)=R\left(\hat{x},\hat{y}\right),\;\;\;t=\frac{R}{V}\hat{t},\\ \mathbf{u}=V\hat{\mathbf{u}},\;\;\;\xi=\frac{V}{R}\hat{\xi},\;\;\;\psi=VR\hat{\psi},\end{array}$ (S3)
@@
$\frac{D}{Dt}\left(\xi+\hat{\beta}y\right)-\frac{1}{{\rm Bu}}\frac{\partial\psi}{\partial t}=0,$ (S4)
@@
$\hat{\beta}\equiv\frac{\beta R^{2}}{V},\;\;\;{\rm Bu}\equiv\frac{L_{d}^{2}}{R^{2}}=\frac{gH}{R^{2}f_{0}^{2}},$ (S5)
@@
$\frac{\partial\psi}{\partial t_{0}}=\frac{\partial\psi}{\partial t}-\hat{\beta}u_{\beta}\frac{\partial\psi}{\partial x_{\beta}},$ (S7)
so the material derivative becomes
$\frac{D_{0}}{Dt}=\frac{\partial_{\beta}}{\partial t}+\left(\mathbf{u}-\hat{\beta}\mathbf{u}_{\beta}\right)\cdot\nabla_{\beta},$ (S8)
[…]
$f=f_{\rm v}(x_{\beta},y_{\beta})+\hat{\beta}f_{\rm g}(x_{\beta},y_{\beta},t).$ (S10)
$\frac{\partial}{\partial t}\left(\xi_{{\rm g}}-\frac{\psi_{{\rm g}}}{{\rm Bu}}\right)=-v_{{\rm v}}-\mathbf{u}_{{\rm v}}\cdot\nabla\xi_{{\rm g}}-\left(\mathbf{u}_{{\rm g}}-\mathbf{u}_{\beta}\right)\cdot\nabla\xi_{{\rm v}}-\frac{1}{{\rm Bu}}\mathbf{u}_{\beta}\cdot\nabla\psi_{{\rm v}}+O(\hat{\beta})=0,$ (S11)
@@
The version of Eq. S11 presented in the main text (Eq. 1) is simplified for focusing the reader’s attention on the important points. It unscales back the "$\rm{g}$" variables by $\hat{\beta}$, so as to directly see the dependence of $\beta$-drift on $\hat{\beta}$, and assumes that the terms $\left(\left(\mathbf{u}_{{\rm g}}-\mathbf{u}_{\beta}\right)\cdot\nabla\xi_{{\rm v}},\;\;-\frac{1}{{\rm Bu}}\mathbf{u}_{\beta}\cdot\nabla\psi_{{\rm v}}\right)$ are relatively small. This assumption can be intuitively understood at the beginning of the motion, where $\mathbf{u}_{{\rm g}}=\mathbf{u}_{\beta}=0$. Nonetheless, the simulations (Fig. 1) use the full version (Eq. S11).
@@
For solving Eq. S11 we use the Dedalus PDE solver Burns et al. (2020) with double periodic bases. To enforce a concentric boundary, we set the (scaled) domain size as $(x,y)\in\{-20,20\}$ and add a relaxation term of the form
$\frac{\xi_{g}}{\tau_{\rm{d}}}G(r,R_{\rm{d}},A_{\rm{d}}),$ (S12)
@@
$G=\frac{1}{2}\left(1-\tanh{\left(A_{\rm{d}}\left(R_{\rm{d}}-r\right)\right)}\right)$ (S13)
@@
$\xi_{\rm v}=\left(2-r^{b}\right)e^{\frac{1}{b}\left(1-r^{b}\right)},$ (S14)
@@
In Fig. 1(a-b), we plot the resulting $(u_{\beta},v_{\beta})$ with $b=1.2$ and four values of $\rm{Bu}$. We find that $u_{\beta}$ behaves like an exponential decay function, $u_{\beta}\sim A\left(1-e^{-t/\tau}\right)$, where the amplitude $A$ and timescale $\tau$ minimize the root-mean-square error (RMSE) relative to the simulated $u_{\beta}$. This parameterization allows for a simple presentation of the behavior of $u_{\beta}$ as a function of $\rm Bu$ (Fig. 1c).
The Polar QG Model
In this section, we present the details of the polar quasi-geostrophic (QG) model and the statistics underlying the results shown in Fig. 2. We solve the SW-QG equation Vallis (2017),
$\frac{D}{Dt}\left(\nabla^{2}\psi+f-\frac{1}{L_{d}^{2}}\psi\right)=0,$ (S15)

### v2

@@ 1 Introduction
Figure 3: Interpretation of the Cyclones’ Vertical Structure from the Deformation Radius. The first row corresponds to the "mode 0" solutions of the eigenvalue problem described by Eq. 6, while the second row corresponds to "mode 1". (a,d) Vertical profiles of $\Phi_{n}$ for each mode. (b,e) Solutions for $L_{d}$ (green shades) for various combinations of NN (Brunt-Väisälä frequency) and $H$ (depth). The estimated $L_{d}$ distributions from Fig. 2k are shown by the blue and orange curves for the north and south poles, respectively. (c,f) Solutions for $\Phi_{n}$ as a function of pressure below the cloud layer, following the NN and $H$ relationship indicated by the blue solid curve in panels (b) and (e). Green horizontal dotted lines indicate the depths at which each MWR channel has maximum sensitivity (Janssen et al., 2017), though each channel is sensitive over a broader range of depths. The numbers on the right ordinate show the altitude relative to 1 bar. Grey dashed lines in panels (b,c,e,f) denote the values used to generate the profiles in panels (a) and (d), and to produce Fig. 4. The color along the grey dashed line in panels (c,f) corresponds to the $\Phi(z)$ profiles shown in panels (a,d). The purple dashed line in panel (c) represents the theoretical scaling $H\sim\mathrm{Ro}^{1/2}Rf_{0}/N$ (adapted from (Aubert et al., 2012)), using representative values of $R=800\,\rm{km}$ and $U=90\,\rm{ms}^{-1}$ from Fig. 2m-n.
@@
Interpreting the Vertical Structure of Jupiter’s Polar Atmosphere from the Drift Rate of the Polar Cyclones
@@
The standard shallow-water (SW) expression for the deformation radius, $L_{d}=\sqrt{gH}/f_{0}$, is not well-suited to Jupiter’s upper atmosphere. This expression assumes a constant density and no vertical dependence of the flow, whereas Jupiter’s upper-atmospheric density increases exponentially with depth Guillot and Morel (1995) over many scale heights. To account for these vertical variations, we consider the continuously stratified QG vorticity equation Vallis (2017):
@@
where $N$ is the Brunt-Väisälä frequency, indicating the stability of the atmosphere with respect to vertical motion, and $\tilde{\rho}$ is the vertical density profile.
@@
where $\Phi_{n}(z)$ is the vertical profile for each $n$. Substituting this form into the QG equations requires each $\Phi_{n}$ to satisfy the vertical eigenvalue problem Smith and Vallis (2001):
@@
where $\Gamma_{n}$ is an eigenvalue. For each mode $n$, the corresponding deformation radius is given by $L_{d,(n)}=\Gamma_{n}^{-\frac{1}{2}}$ (see SI for the full derivation). Assuming that the cyclones are dominated by a single vertical mode, the horizontal dynamics derived from the two-dimensional QG equation (Eq. 3), and the three-dimensional QG equation (Eq. 4) will yield similar results for the given $L_{d,(n)}$; the linearized dynamics are identical (see SI). Given our estimates of the $L_{d}$ values required to achieve the observed westward drift (Fig. 2), we can use the eigenvalue problem posed by Eq. 6 to investigate which vertical structures (eigenfunctions $\Phi_{n}(z)$) satisfy the required $L_{d}$.
@@
We solve this eigenvalue problem using the Dedalus library Burns et al. (2020), applying boundary conditions of no vertical velocity at the tropopause $(\partial_{z}\Phi_{n}|_{{}_{z=0}}=0)$ and no flow at the bottom $(\Phi_{n}|_{{}_{z=-H}}=0)$. For $\tilde{\rho}(z)$, we adopt a reference density profile derived from interior models of Jupiter, constrained by a neural-network-based parameter sweep Ziv et al. (2024a, b). We focus here on the two lowest modes ($n=\{0,1\}$), which are illustrated in Fig. 3(a,d). The $n=0$ mode corresponds to a simple decay profile, whereas $n=1$ changes sign along the column. This focus on the two leading modes aligns with estimates from Earth’s oceans, where they capture the bulk of the kinetic energy Smith and Vallis (2001); Ferrari and Wunsch (2010); higher modes require excitations with elaborate vertical profiles.
@@
Since Eq. 6 depends only on $H$ and $N$ (here assumed constant, see discussion in the SI), we present $L_{d,(n)}$ solutions as a function of these two parameters in Fig. 3(b,e), for both modes. The constrained range of deformation radii from the previous section (Fig. 2k) is indicated for the north (blue) and south (orange) poles, essentially giving an $H$ as a function of $N$ for each pole. Focusing on the relationship defined by the blue solid curve in Fig. 3(b,e), we plot the corresponding $\Phi_{n}(z)$ profiles as a function of pressure below the cloud level for different $N$ values (Fig. 3(c,f)).
Interestingly, in agreement with the eigenvalue results, the $H$ vs $N$ line in Fig. 3c closely follows a scaling law derived for the aspect ratio of vortices in stratified fluids, supported by theory and rotating tank experiments (Billant and Chomaz, 2001; Aubert et al., 2012; Lemasquerier et al., 2020). This vertical scaling, $H\sim f_{0}R\sqrt{\mathrm{Ro}\left(1-\mathrm{Ro}\right)/\left(N_{c}^{2}-N^{2}\right)}$, can be simplified in our case to $H\sim f_{0}R\,\sqrt{\mathrm{Ro}}/N$. This approximation is valid under the assumptions that the Rossby number is small ($\mathrm{Ro}=U/(f_{0}R)\ll 1$) and that the static stability at the vortex center is small ($N_{c}^{2}\ll N^{2}$). Under these conditions, the simplified scaling (purple dashed line in Fig. 3c) provides a good match to the curve derived from the eigenvalue solutions for the North Pole.
@@
The MWR instrument, expected to probe below Jupiter’s clouds at the north pole in the near future, measures brightness temperature at six channels. The maximum sensitivity pressure levels for each MWR channel are marked by green dotted lines, with Channel 6 above the 1-bar level. For a plausible range $N=3-20\times 10^{-3}\rm{s}^{-1}$ Lee and Kaspi (2021), consistent with our reanalysis of Galileo probe data (see SI and Fig. S5), and focusing first on mode 0 (Fig. 3c), we anticipate that the north polar cyclones’ footprint will be visible in Channels 6 and 5 (and Channel 4 if $N\approx 3\times 10^{-3}\rm{s}^{-1}$). However, if the real Jovian poles exhibit turbulence with a vertical phase shift, reminiscent of terrestrial storms with low-level convergence and upper-level divergence, then mode 1 may also be excited. In that scenario (Fig. 3f), signatures could extend into Channel 3 for the same range of $N$s.
@@
Figure 4 illustrates an example of how $\psi$ extends downward and intersects with the MWR channels for $N=3\times 10^{-3}\,\mathrm{s}^{-1}$, the lower bound estimated in Lee and Kaspi (2021). Channel 6, sensitive to pressures around $0.7$ bar, lies slightly above the eigenfunctions’ range (which starts near 1 bar). We assume that Channel 6 will display a $\psi$ field similar to that of Channel 5. The JIRAM background photograph Adriani et al. (2018) represents the cloud deck at approximately this pressure level. The downward extension of $\psi$ below Channel 4 depends on the excitation of mode 1. It is important to note that the MWR measures brightness temperature, which is only indirectly related to $\psi$. Predicting MWR observations requires a layer of translation, involving solutions for vertical velocity and the subsequent diffusion-advection of ammonia, given the background ammonia distribution Duer et al. (2021).
@@
Figure 4: Vertical Structure and Intersections with Instrument Sensitivity Depths at the North Pole. The $\psi$ field from Fig. 2c was extended downward (rather than computed directly in a 3D simulation) using the eigenfunctions from Fig. 3a,d, which correspond to $N=3\times 10^{-3}\;(\mathrm{s}^{-1})$ and $H=50$ ($133$) km for mode 0 (mode 1). Both modes yield $L_{d}=226$ km, with mode 0 terminating ($\psi=0$) at z=50z=-50 km, and mode 1 continuing further to $\psi=0$ at z=134z=-134 km. The planes represent $\psi$ at the MWR channels that overlap with the eigenfunction range, while the shaded regions between the planes depict $\psi$ values between the channels. The dashed black line indicates the north pole. The background image on channel 6 is a JIRAM measurement from PJ4 Adriani et al. (2018), similar to Fig. 2a.
Discussion
@@
We then employed a 2D polar model initialized with identical cyclones at their observed locations (during PJ4). This model provides a dynamical constraint on the deformation radius, $L_{d}$ (Fig. 2), which encodes information about the cyclones’ vertical structure. We found that both excessively large and excessively small values of $L_{d}$ produce westward drifts that deviate strongly from observations. A sampling algorithm estimates $L_{d}\sim 226\,(\mathrm{km})$ at the north pole and $\sim 364\,(\mathrm{km})$ at the south pole. While the posterior distributions of $L_{d}$ at the two poles do overlap, it remains an open question whether genuine differences in $L_{d}$ exist between the poles, especially given Jupiter’s negligible obliquity and the resulting expected hemispheric symmetry. However, the indicated difference in $L_{d}$ between the poles could have important implications for the characteristics of the polar cyclones. Since $L_{d}$ controls the spatial scale of vortices in geostrophic turbulence (Polvani et al., 1994), the possibly smaller $L_{d}$ at the north pole may explain the greater number of smaller cyclones observed at the north pole compared to the south.
Using these estimates of $L_{d}$, we sought consistency between the 2D QG framework (Fig. 2) and a 3D QG model that incorporates density variations with depth (Eq. 4). This required solving the eigenvalue problem in Eq. 6, in which the eigenvalues yield $L_{d}$ for different vertical modes, and the numerically determined eigenfunctions describe the cyclone’s vertical structure relative to the cloud deck. In our idealized approach, we assumed a constant Brunt-Väisälä frequency, $N$, and a single dominant vertical mode, both of which can introduce uncertainty.
We emphasize that the eigenvalue-derived $L_{d}$ represents a vertically integrated quantity that encapsulates, as accurately as possible, the effects of vertical stretching on 2D dynamics when density varies significantly with depth and the atmosphere is stably stratified. While the linear dynamics at the top layer ($z=0$) of the 3D-QG model (Eq. 4) and the 2D-QG model (Eq. 3) are identical—for example, in a Rossby wave solution Vallis (2017)—nonlinearities and multi-modal dynamics introduce deviations between the models (see SI). Quantifying this error is worthwhile but requires full 3D simulations at the pole with sufficient resolution, which is left for future work.
The solutions (Fig. 3) suggest that for $N\approx 10^{-2}\,\mathrm{s}^{-1}$, cyclone depths of 107510-75\, km at the north pole and 209020-90\, km at the south pole are plausible, depending on the mode. If $N\approx 3\times 10^{-3}\,\mathrm{s}^{-1}$, the estimated cyclone depths increase substantially, ranging 3022030-220 km at the north pole and 6033060-330\, km at the south pole. As a comparison, the GRS has been estimated to be approximately $300\pm 100$ km deep based on gravity measurements Parisi et al. (2021), and its signature has been observed down to MWR channel 1, corresponding to depths of $O(300\,\mathrm{km}$) Bolton et al. (2021), consistent with both experimental and numerical modeling results Lemasquerier et al. (2020). Although the GRS is a distinct low-latitude anticyclone, these estimates suggest that the lower-N ($\sim 3\times 10^{-3}\,\rm{s}^{-1}$) regime considered here may provide a more realistic representation of Jupiter’s deep vortices.
@@
In summary, our work presents a novel framework connecting observed cloud-level motions to the deeper structure of Jovian polar cyclones. By linking the deformation radius, derived from $\beta$-drift constraints, to vertical modes in a 3D-QG framework, we gain new insights into how these cyclones extend below the visible cloud layer and how their internal structure influences their horizontal drift. These constraints on deformation radius, static stability, and vertical profile can guide future studies aimed at modeling the energy fluxes responsible for the formation and maintenance of these cyclones against dissipation.
@@
No new data sets were generated during the current study. The data analyzed in this study were published by Ref.Mura et al. (2022) (DOI: https://doi.org/10.1029/2022JE007241), as cited in the text. The simulations were run with the Dedalus solver Burns et al. (2020) (DOI: https://doi.org/10.1103/PhysRevResearch.2.023068).
@@ Acknowledgments
This research has been supported by the Minerva Foundation with funding from the Federal German Ministry for Education and Research and the Helen Kimmel Center for Planetary Science at the Weizmann Institute of Science. We thank Rei Chemke, Keren Duer, Eli Galanti and Or Hadas for insightful conversations and helpful feedback.
@@
Appendix A Developing a Model Following a Cyclone That Includes the Effect of Stretching
@@
$\frac{D}{Dt}\left(\xi+\beta y-\frac{1}{L_{d}^{2}}\psi\right)=0,$ (9)
@@
$\frac{D}{Dt}=\left(\frac{\partial}{\partial t}+\mathbf{u}\cdot\nabla\right),$ (10)
$\mathbf{u}$ is the velocity vector, $\xi=\nabla\times\mathbf{u}$ is the relative vorticity, $\beta$ is the constant meridional gradient of the planetary vorticity ($f$) or any prescribed background vorticity gradient, $L_{d}=\sqrt{gH}/f_{0}$ is the deformation radius, $g$ is the gravitational acceleration, $H$ is the layer’s depth, and $f_{0}$ is the reference background rotation rate.
@@
[…]
@@
$\frac{D}{Dt}\left(\xi+\hat{\beta}y\right)-\frac{1}{{\rm Bu}}\frac{\partial\psi}{\partial t}=0,$ (12)
@@
$\hat{\beta}\equiv\frac{\beta R^{2}}{V},\;\;\;{\rm Bu}\equiv\frac{L_{d}^{2}}{R^{2}}=\frac{gH}{R^{2}f_{0}^{2}},$ (13)
@@
$\begin{array}[]{c}x=x_{\beta}+\int_{0}^{t}\hat{\beta}u_{\beta}dt,\\ y=y_{\beta}+\int_{0}^{t}\hat{\beta}v_{\beta}dt,\end{array}$ (14)
$\frac{\partial\psi}{\partial t_{0}}=\frac{\partial\psi}{\partial t}-\hat{\beta}u_{\beta}\frac{\partial\psi}{\partial x_{\beta}}-\hat{\beta}v_{\beta}\frac{\partial\psi}{\partial y_{\beta}},$ (15)
so the material derivative becomes
$\frac{D_{0}}{Dt}=\frac{\partial_{\beta}}{\partial t}+\left(\mathbf{u}-\hat{\beta}\mathbf{u}_{\beta}\right)\cdot\nabla_{\beta},$ (16)
$\biggl[\frac{\partial}{\partial t}+\left(\mathbf{u}-\hat{\beta}\,\mathbf{u}_{\beta}\right)\cdot\nabla_{\beta}\biggr]\Bigl(\,\xi+\hat{\beta}\bigl[y+\hat{\beta}\!\int_{0}^{t}v_{\beta}\,dt\bigr]\Bigr)\;-\;\frac{1}{\mathrm{Bu}}\biggl(\frac{\partial\psi}{\partial t}-\hat{\beta}\,\mathbf{u}_{\beta}\cdot\nabla_{\beta}\psi\biggr)\;=\;0.$ (17)
$f=f_{\rm v}(x_{\beta},y_{\beta})+\hat{\beta}f_{\rm g}(x_{\beta},y_{\beta},t).$ (18)
@@
The version of Eq. 19 presented in the main text (Eq. 1) is simplified for focusing the reader’s attention on the important points. It unscales back the $\rm{g}$" variables by $\hat{\beta}$, so as to directly see the dependence of $\beta$-drift on $\hat{\beta}$, and assumes that the terms $\left(\left(\mathbf{u}_{{\rm g}}-\mathbf{u}_{\beta}\right)\cdot\nabla\xi_{{\rm v}},\;\;-\frac{1}{{\rm Bu}}\mathbf{u}_{\beta}\cdot\nabla\psi_{{\rm v}}\right)$ are relatively small. This assumption can be intuitively understood at the beginning of the motion, where $\mathbf{u}_{{\rm g}}=\mathbf{u}_{\beta}=0$. Nonetheless, the simulations (Fig. 1) use the full version (Eq. 19).
@@
For solving Eq. 19 we use the Dedalus PDE solver Burns et al. (2020) with double periodic bases. To enforce a concentric boundary, we set the (scaled) domain size as $(x,y)\in\{-20,20\}$ and add a relaxation term of the form
$\frac{\xi_{g}}{\tau_{\rm{d}}}G(r,R_{\rm{d}},A_{\rm{d}}),$ (20)
@@
$G=\frac{1}{2}\left(1-\tanh{\left(A_{\rm{d}}\left(R_{\rm{d}}-r\right)\right)}\right)$ (21)
@@
$\xi_{\rm v}=\left(2-r^{b}\right)e^{\frac{1}{b}\left(1-r^{b}\right)},$ (22)
@@
Figure S2: Analysis of the exponential fit, $\rm{U}_{\beta}(1-e^{-t/\tau})$, adequacy. (a) The same as Fig. 1a, but for different $\rm Bu$ values, showing poor exponential fit for $\rm{Bu}=(7.9,8.89)$. (b) The mean integrated absolute error between the $u_{\beta}$ and the exponential fit. It can be seen that for $\rm Bu\gtrsim 5$ the exponential parametrization is not adequate for capturing $u_{\beta}(t)$.
In Fig. 1(a-b), we plot the resulting $(u_{\beta},v_{\beta})$ with $b=1.2$ and four values of $\rm{Bu}$. We find that $u_{\beta}$ behaves like an exponential relaxation function, $u_{\beta}\sim\rm{U}_{\beta}\left(1-e^{-t/\tau}\right)$, where the amplitude $\rm{U}_{\beta}$ and timescale $\tau$ minimize the root-mean-square error (RMSE) relative to the simulated $u_{\beta}$. This parameterization allows for a simple presentation of the behavior of $u_{\beta}$ as a function of $\rm Bu$ (Fig. 1c). We observe that this parameterization becomes poorly fitted for values of $\rm Bu\gtrsim 5$, as illustrated in Fig. S2.
As seen in Fig. 1a–b, $\mathbf{u}_{\beta}$ exhibits oscillations at the beginning of the motion. This likely results from the following mechanism: the vortex core rotates significantly faster than its outer regions, and $\beta$-gyres initially develop within the core, aligned in the east–west direction. These gyres begin to rotate with the core, generating the observed wiggling in the $\beta$-drift velocity. The oscillations gradually subside as the core homogenizes the background vorticity gradient and as the gyres propagate outward from the core. This behavior is illustrated in Movie S1, which shows the simulation corresponding to Fig. 1e. A Dedalus model script for solving Eqs. 1922 is available on GitHub (beta_plane_moving_frame_vortex.py).
Appendix B The Polar QG Model
In this section, we present the details of the polar QG model and the statistics underlying the results shown in Fig. 2. We solve the 2D-QG equation Vallis (2017),
$\frac{D}{Dt}\left(\nabla^{2}\psi+f-\frac{1}{L_{d}^{2}}\psi\right)=0,$ (23)